# The Study of the Canonical forms of Killing tensor in vacuum with $\Lambda$

### v1

@@ Abstract
This paper is the initial part of a comprehensive study of spacetimes that admit the canonical forms of Killing tensor in General Relativity. Our scope is to derive either new exact solutions of Einstein’s equations or to determine the hidden symmetries of the already known ones. In this preliminary work we first introduce the canonical forms of Killing tensor. Subsequently, we employ the integrability conditions of each canonical form along with the Einstein field equations (in vacuum with $\Lambda$) and the Bianchi identities in an attempt to create a solvable yet overdetermined system of equations. Finally, we obtain multiple special algebraic solutions according to the Petrov classification (D, III, N, O). The latter becomes possible since our analysis is embodied with the usage of the Newman-Penrose formalism of null tetrads.
Keywords: Canonical forms, Killing tensor, Newman-Penrose formalism, Exact solutions,
Einstein’s equations, Type D, Type III, Type N
@@ 1 Introduction
The interpretation and the analysis of exact solutions in the context of General Relativity constitutes a whole regime of research. It is already within our understanding that the Einstein’s Field Equations (EFE) cannot be solved without additional mathematical assumptions regarding the nature of a spacetime. Thus, the consideration of symmetries in the resolution process of the EFE is indispensable. In other words, the non-linear character of the equations obligate us to introduce further information through symmetries to obtain a solvable-overdetermined system. In this work we are not interested in any kind of symmetry, but we put under the spotlight spacetimes with hidden symmetries. We aim to find spacetimes which admit Killing Tensors since they are responsible for both explicit and hidden symmetries.
A well known example of a hidden symmetry is the conservation of Laplace-Runge-Lenz vector in the Kepler-Coulomb problem along geodesics as a constant of motion which is connected with the existence of Killing-Yano Tensor [Papakostas, 2001], a generalization of Killing Tensor [Cariglia, 2014],[Papakostas, 1985]. Also, a relevant conjecture says that: in a spacetime, which admit a non-trivial Killing tensor [^1] The trivial Killing tensor is the metric tensor $g_{\mu\nu}$ where its existence indicates the conservation of the rest mass of a moving particle in Hamiltonian systems., there are closed trajectories [Burns and Matveev, 2021].
Furthermore, as Eisenhart [Eisenhart, 1934] and Kalnins-Miller [Kalnins and Miller, 1980], [Kalnins and Miller, 1981], [Kalnins and Miller, 1983] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of valence two [Benenti, 2016]. Indeed, there is a bizarre relation between the structure of separated metrics with the structure of its characteristic Killing tensor. Benenti and Francaviglia [Benenti and Francaviglia, 1979] presented a certain example where the additional information about the metric tensor serve as a catalyst in order to obtain the structure of the Killing tensor.
Besides, there are two ways to benefit from Killing tensor, either, by assuming its existence, to find a metric or by revealing the hidden symmetries of a known metric or both. Consequently, the existence of Killing tensor in a physical problem helps us to determine a solvable system of equations through its integrability conditions (IC) and, in some cases, leads to the separation of Hamilton-Jacobi equation providing us with integrable trajectories. Regarding this, the assumption of existence of Killing tensor could serve as a promising starting point in the pursuit of “realistic” spacetimes endowed by integrable trajectories.
@@
The main idea that motivated us to deal with the Canonical Forms of the Killing tensor was that we might find new interesting spacetimes in vacuum if we deal with more general forms of a Killing tensor with more than two distinct eigenvalues as postulated by Hauser and Malhiot in [Hauser and Malhiot, 1976]. During the last decades the only works in the literature that utilize a Killing tensor to explore new spacetimes including Hauser and Malhiot’s work on electro-vacuum [Hauser and Malhiot, 1978] and the work of one of us on interior solutions with perfect fluid [Papakostas, 1998]. Hauser-Malhiot managed to found one of the most general family of stationary axially symmetric electro-vacuum spacetimes that found independently by Carter [Carter, 1968]. But both of these works serve as the only paradigms for our research, since these are the only works both originate from the assumption of existence of a Killing tensor with two double eigenvalues.
It is worth mentioning that the Killing tensor form with two double eigenvalues is a special case of the canonical forms of a Killing tensor, and its study has yielded general and new families of exact solutions in electro-vacuum.
@@
It is both interesting and important to explore Einstein’s spacetimes that admit more general forms of Killing tensors beyond the case of two double eigenvalues as initial premise. These more general forms represent the four canonical forms of Killing tensor. It is also intriguing to investigate whether these more general forms can lead to new solutions or generalizations of the ones already known. Additionally, we aim to investigate whether more general Petrov type solutions (e.g. Petrov type I) admit the canonical forms of the Killing tensor, as opposed to the Killing form of the Paradigm. We choose thus to take advantage of the canonical forms of Killing Tensor to discover new metrics with explicit and hidden symmetries or to acquire more information of already known metrics.
It’s also worth noting that the use of the standard metric formalism is not practical for our purpose, as it cannot accommodate the symmetries associated with a Killing tensor. In this context, the most suitable formalism is the complex vectorial formalism of spin coefficients [Cahen et al., 1967] or the Newman-Penrose formalism [Newman and Penrose, 1962]. The first reason is that during the resolution process, the classification of the gravitational field of a spacetime according to Petrov occurs in the early stages of the process. This allows us to deduce theorems of symmetries for each Petrov type [Petrov, 2000], determining an appropriate coordinate system using the Frobenius theorem of integrability.
@@
We have attempted to establish a coherent structure for this work. In Chapter 2 we will exhibit the main points of the appointed formalism which will be revisited throughout the paper. Next, in Chapter 3 we imply a rotation around the null tetrad frame with $l^{\mu}$ fixed and we obtain the general key relations that can be applied to any canonical form except $K^{0}$. In Chapter 4 the definition of Killing tensor is given and we revisit the canonical forms of Killing tensor. After that in Chapters 5, 6 and 7 the resolution process takes place for $K^{0}$, $K^{1}$ and $K^{2,3}$, therein we present the Petrov types of the obtained solutions. The Chapter 8 before Discussion and Conclusions contains a reduced Killing form with $\lambda_{0}=0$ which is a subcase of $K^{1},K^{2},K^{3}$ Killing forms.
@@ 2 Notation of the Newman-Penrose Formalism
The Newman-Penrose Formalism is a widely known formalism that was presented by Newman and Penrose [Newman and Penrose, 1962] and was analyzed geometrically by Cahen, Debever and Defrise [Cahen et al., 1967], [Debever, 1964]. Initially, the formalism was found in order to describe the gravitational radiation in General Relativity but it was proved to have much more usefulness.
@@ 2.2 Bivector Space
The Ricci tensor components are represented by $E_{a\bar{b}}$ and they are divided in to real and complex components. All these quantities describe the main parts of the EFE. The EFE in this formalism are represented by the corresponding field equations which are known either as Newman-Penrose Equations (NPE) or as Ricci identities [Newman and Penrose, 1962].
@@
All the above sets of equations contribute to the NPE, the BI and the CR of the basis vectors. Despite the fact that we have to solve a considerably larger number of equations this formalism has great advantages. Gauge transformations of the tetrad can be used to simplify the field equations and we can easily extract invariant properties of the gravitational field (Petrov types) without using a coordinate basis [Stephani et al., 2009]. Also, it allows us to search for solutions with specific special features, such as the presence of one or two null directions that might be singled out by physical or geometric considerations.
@@ 3 Null rotation
In this work we choose to take advantage of the conformal symmetry of a rotation around one of the real null vectors, namely we choose $l^{\mu}$ to be fixed, although we will exhibit both transformations in the relations below, wherein the exponentials correspond to Lorentz transformations and the parentheses correspond to null rotation [Stewart, 1993]. We define the complex rotation parameters $t\equiv a+ib$ and $p\equiv c+id$.
@@
There are two different kinds of simplifications that can be acquired by the capitalization of the annihilation of the tilded spin coefficients $\tilde{\epsilon},\tilde{\gamma},\tilde{\alpha},\tilde{\beta}$. The simplest simplification emerges by the correlation of the spin coefficient with the derivative of the rotation parameter $t$. In case where $\lambda_{0}=0$ the rotation parameter becomes $t=a$, differently, in case where $\lambda_{7}$ the non-zero rotation parameter is $t=ib$. The latter has significant impact on the spin coefficients. When $t=a$ the spin coefficients $\epsilon,\gamma,\alpha,\beta$ depends on the directional derivatives of real quantities yielding the following relations
@@
The second kind of simplification takes place when we substitute the four spin relations into the CR and we compare the outcome with the NPE (f),(l),(e),(o) resulting in the key relations. The latter unfolds the branches of the solutions. We postulate the most general case of the obtained relations after the comparison with NPE.
@@
Remark The application of null rotation and Lorentz transformations are transformative processes that yield valuable relationships connecting not only the spin coefficients amongst themselves but also with the Weyl components through the commutation relations, once the tilded spin coefficients have been annihilated. The outcome critically hinges on the preservation of the structure of the Killing tensor and the form of the spin coefficients. Importantly, the absence of either $\lambda_{0}$ or $\lambda_{7}$ is the catalyst for these relationships. Employing the complete form of the Killing tensor $K^{1,2,3}_{\mu\nu}$, conversely, provides no insight into this implied symmetry since both $t$ and $p$ are nullified.
@@ 4 The canonical forms of Killing tensor
The preservation of geometry of spacetime through a transformation reveals the existence of symmetries. These symmetries leave invariant the elements which characterize the geometry, that is referred to the metric tensor and, by extension, the action. Thus, the consideration of symmetries in the resolution process of the Einstein’s equations is indispensable. The non-linear character of the equations of gravity obligates us to introduce further information through symmetries to obtain a solvable-overdetermined system of equations.
@@ 4.1 Killing Tensor
where $p_{\mu}$ are the coordinates on the cotangent spaces or equivalently the canonical momenta of an observer. Then, an integral of motion could be defined as follows [Kruglikov and Matveev, 2016], [Sommers, 1973].
@@
Killing tensors of rank r give rise to a homogeneous constant of motion of degree r in momenta. The inhomogeneous polynomial integrals of geodesic motion can be decomposed to their homogeneous parts and also to the corresponding parts that are associated with the Killing tensors with the equivalent rank. The Killing tensor of rank 1 equals to Killing vector and generates continuous symmetry transformations, these symmetries are called explicit [Garfinkle and Glass, 2010]. The symmetries that correspond to higher-order ranks of momenta associated with Killing tensors of rank ($r>1$) are called hidden symmetries [Frolov et al., 2017]. The investigation of special objects such Killing tensor or Killing-Yano is basically a devilish way to peep into the phase space searching for hidden symmetries [Krtouš et al., 2007]. Indeed, the explicit symmetries in a Hamiltonian system always could be “dragged up” instead of hidden symmetries. These kinds of symmetries emerge during the study of the dynamics of a system featuring the conserved quantities of the system or one-parameter isometries which is equivalent with the admission of existence of Killing vectors.
As Eisenhart [Eisenhart, 1934] and Kalnins-Miller [Kalnins and Miller, 1980], [Kalnins and Miller, 1981], [Kalnins and Miller, 1983] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of order two [Benenti, 2016]. Indeed, there is a bizarre relation between the structure of separated metrics with the structure of its characteristic Killing tensor. Benenti and Francaviglia [Benenti and Francaviglia, 1979] present a certain example where the additional information about the metric tensor serve as a catalyst in order to obtain the structure of the Killing tensor.
Let us proceed with the definition of Killing tensor: Any symmetric tensor of order 2 whose the symmetric part of his covariant derivative vanishes is called a Killing tensor.

### v2

@@ Abstract
This paper is the initial part of a comprehensive study of spacetimes that admit the canonical forms of Killing tensor in General Relativity. The general scope of the study is to derive either new exact solutions of Einstein’s equations that exhibit hidden symmetries or to identify the hidden symmetries in already known spacetimes that may emerge during the resolution process. In this preliminary paper, we first introduce the canonical forms of Killing tensor, based on a geometrical approach to classify the canonical forms of symmetric 2-rank tensors, as postulated by R. V. Churchill. Subsequently, the derived integrability conditions of the canonical forms serve as additional equations transforming the under-determined system of equations, comprising of Einstein’s Field Equations and the Bianchi Identities (in vacuum with $\Lambda$), into an over-determined one. Using a null rotation around the null tetrad frame we manage to simplify the system of equations to the point where the geometric characterization (Petrov Classification) of the extracted solutions can be performed and their null congruences can be characterized geometrically. Therein, we obtain multiple special algebraic solutions according to the Petrov classification (D, III, N, O) where some of them appeared to be new. The latter becomes possible since our analysis is embodied with the usage of the Newman-Penrose formalism of null tetrads.
keywords:
Canonical forms, Killing tensor, Newman-Penrose formalism, Exact solutions, Einstein’s equations, Type D, Type III, Type N
@@ 1 Introduction
The objective of this paper is to find spacetimes with hidden symmetries, which are exact solutions of Einstein’s equations in vacuum with cosmological constant $\Lambda$, using the existence of an abstract 2-rank Killing tensor as the only premise. It should be noted that no additional assumptions are made regarding specific Petrov types, invertibility, separability, or other conditions. For reasons of clarity, we begin with a review in which our motives, goals and outcomes will be further explained, providing the reader with a broader understanding of the scope of this work.
The interpretation and the analysis of exact solutions in the context of General Relativity constitute a whole regime of research. In this scientific branch exact solutions represent the hidden trophy within the non-linear nature of the equations of gravity. These solutions can be obtained by imposing restrictions on the under-determined system of equations consisting of Einstein’s Field Equations (EFE) and Bianchi Identities (BI) [^1] Since this work was operated in the framework of Newman-Penrose formalism, the EFE corresponds to the Newman-Penrose Equations (NPE).. Such restrictions may be related to the algebraic structure of the Riemann tensor, to symmetry conditions on the metric, to the field equations of the stress energy-momentum tensor and other related considerations.
The appropriate way to obtain exact solutions based on the bespoken assumption is to solve a system of equations regarding the EFE, BI and the Integrability Conditions (IC) of each canonical form of Killing tensor. According to this, we begin by introducing the canonical forms of the Killing tensor and deriving the integrability conditions for each of them. Hence, the resolution process initiates by constructing an over-determined-solvable system of equations consisting of EFE, BI, and the IC of each Killing form. The latter description makes clear that the canonical forms is the cornerstone of this problem.
The derivation of the canonical forms of a symmetric matrix will be proved a quite challenging task if someone attempts to do it algebraically. Churchill in 1932 [1] presented the canonical forms of a symmetric 2-rank tensor in pseudo-Euclidean spacetime based on a geometrical approach. This geometrical approach is based on 1) a theorem of Rainich regarding the existence of invariable planes within spacetime [2] and 2) the existence of null vectors within these planes. We managed to acquire the canonical forms of Killing tensor in the framework of Newman-Penrose formalism [3] following the exact same procedure as that operated by Churchill, considering an opposite signature though (Chapter 4).
The next step in the resolution process is to determine the system of equations (EFE, BI, IC) and to operate a null rotation around the null tetrad frame wherein the Killing form must remains invariant. This transformation was used to constrain the arbitrariness of the null tetrad frame while simultaneously simplifying the system of equations.
In this preliminary work, the resolution procedure stopped at the point where the classification of the gravitational field according to Petrov [4] and the characterization of the null congruences take place resulting in to six theorems. We managed to extract all the non-conformally flat solutions in vacuum with $\Lambda$ using as an assumption the existence of canonical forms of Killing tensor. As a matter of fact, in Chapters 7, 8 the type D solutions with a shear-free and non-geodesic null congruence and with a shearing and geodesic null congruence that described by Theorems 4, 5 accordingly, appear to be new as we claim in Chapter 9. We intuit that they could be possible reductions (in vacuum with $\Lambda$) of the undiscovered solutions with Principal Null Directions of Einstein-Maxwell tensor aligned (or non-aligned) with the Principal Null Directions of Weyl tensor [5]. Apparently, this work paves the way to find their spacetime metrics in full detail by solving the remaining equations, but first determining a coordinate system using the Frobenius theorem of integrability.
@@
It should be noted that the main idea that motivated us to deal with the canonical forms of Killing tensor was that we might find new interesting spacetimes in vacuum if we deal with more general forms of a Killing tensor with more than two distinct eigenvalues as postulated by Hauser and Malhiot in [6]. During the last decades the only works in the literature that utilize a Killing tensor to explore new spacetimes including Hauser and Malhiot’s work on electro-vacuum [7] and the work of one of us on interior solutions with perfect fluid [8]. Hauser-Malhiot managed to found one of the most general family of stationary axially symmetric electro-vacuum spacetimes that found independently by Carter [9]. But both of these works serve as the only paradigms for our research, since these are the only works both originate from the assumption of existence of a Killing tensor with two double eigenvalues only. It is worth mentioning that the Killing tensor form with two double eigenvalues is a special case of the canonical forms of a Killing tensor, and its study has yielded general and new families of exact solutions in electro-vacuum.
@@
It is both interesting and important to explore Einstein’s spacetimes that admit more general forms of Killing tensors beyond the case of two double eigenvalues as initial premise. It is also intriguing to investigate whether the canonical forms can lead to new solutions or generalizations of the ones already known. Additionally, in the future we aim to investigate whether more general Petrov type solutions (e.g. Petrov type I) admit the canonical forms of the Killing tensor, as opposed to the Killing form of the Paradigm.
We are intrigued to study spacetimes with hidden symmetries derived from Killing tensors, because there are indications related to the physical nature of the trajectories within these spacetimes. It is crucial to consider that the observed trajectories are closed. A relevant conjecture suggests that in a spacetime admitting a non-trivial Killing tensor [^2] The trivial Killing tensor is the metric tensor $g_{\mu\nu}$ where its existence indicates the conservation of the rest mass of a moving particle in Hamiltonian systems., closed trajectories are present[10].
Furthermore, as Eisenhart [11] and Kalnins-Miller [12], [13], [14] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of valence two [15] in some cases, leads to the separation of Hamilton-Jacobi equation providing us with integrable trajectories. Regarding this, the assumption of existence of Killing tensor could serve as a promising starting point in the pursuit of “realistic” spacetimes endowed by integrable trajectories.
For instance, a well known example of a hidden symmetry is the conservation of Laplace-Runge-Lenz vector in the Kepler-Coulomb problem along geodesics as a constant of motion which is connected with the existence of Killing-Yano Tensor [16], a generalization of Killing Tensor [17],[18]. Another well known example of this kind of symmetry is the Carter’s constant which is associated with probe inclination [19] and it is considered the fourth constant of motion in Kerr-like geometry.
Consequently, entangling the Killing tensor in a physical problem we achieve a twofold result: on one hand, by capitalizing on the Killing tensor, we incorporate additional equations (IC) into the previously under-determined system (EFE, BI), thereby transforming it into an over-determined one; on the other hand, the resulting spacetimes are endowed with hidden symmetries.
It’s also worth noting that the use of the standard metric formalism is not practical for our purpose, as it cannot accommodate the symmetries associated with a Killing tensor. In this context, the most suitable formalism is the complex vectorial formalism of spin coefficients [20] or the Newman-Penrose formalism [21]. The first reason is that during the resolution process, the classification of the gravitational field of a spacetime according to Petrov occurs in the early stages of the process. This allows us to deduce theorems of symmetries for each Petrov type, determining an appropriate coordinate system using the Frobenius theorem.
@@
We have attempted to establish a coherent structure for this work. In Chapter 2 we will exhibit the main points of the appointed formalism which will be revisited throughout the paper. Next, in Chapter 3 we imply a rotation around the null tetrad frame with $l^{\mu}$ fixed and we obtain the general key relations that can be applied to any canonical form except $K^{0}$. In Chapter 4 the definition of Killing tensor is given and we introduce the canonical forms of Killing tensor in the framework of Newman-Penrose formalism based on Churchill’s work . After that, in Chapters 5, 6 and 7 the resolution process takes place for $K^{0}$, $K^{1}$ and $K^{2,3}$, therein we present the Petrov types of the obtained solutions and we determine the characteristics of their null congruences. The Chapter 8 before Discussion and Conclusions contains a reduced Killing form with $\lambda_{0}=0$ which is a subcase of $K^{1},K^{2},K^{3}$ Killing forms. In the final chapter of this work (Chapter 9), we briefly discuss, among other things, the role of entangled transformations in the resolution process (Lorentz transformations and null rotations) with the aim of obtaining exact solutions for more general Petrov types.
@@ 2 Notation of the Newman-Penrose Formalism
The Newman-Penrose Formalism is a widely known formalism that was presented by Newman and Penrose [21] and was analyzed geometrically by Cahen, Debever and Defrise [20], [22]. Initially, the formalism was found in order to describe the gravitational radiation in General Relativity but it was proved to have much more usefulness.
@@ 2.2 Bivector Space
The Ricci tensor components are represented by $E_{a\bar{b}}$ and they are divided in to real and complex components. All these quantities describe the main parts of the EFE. The EFE in this formalism are represented by the corresponding field equations which are known either as Newman-Penrose Equations (NPE) or as Ricci identities [21].
@@
All the above sets of equations contribute to the NPE, the BI and the CR of the basis vectors. Despite the fact that we have to solve a considerably larger number of equations this formalism has great advantages. Gauge transformations of the tetrad can be used to simplify the field equations and we can easily extract invariant properties of the gravitational field (Petrov types) without using a coordinate basis [23]. Also, it allows us to search for solutions with specific special features, such as the presence of one or two null directions that might be singled out by physical or geometric considerations.
@@ 3 Null rotation
In this work we choose to take advantage of the conformal symmetry of a rotation around one of the real null vectors, namely we choose $l^{\mu}$ to be fixed, although we will exhibit both transformations in the relations below, wherein the exponentials correspond to Lorentz transformations and the parentheses correspond to null rotation [24]. We define the complex rotation parameters $t\equiv a+ib$ and $p\equiv c+id$.
@@
There are two different kinds of simplifications that can be acquired by the capitalization of the annihilation of the tilded spin coefficients $\tilde{\epsilon},\tilde{\gamma},\tilde{\alpha},\tilde{\beta}$. The simplest simplification emerges by the correlation of the spin coefficient with the derivative of the rotation parameter $t$. In case where $\lambda_{0}=0$ the rotation parameter becomes $t=a$, differently, in case where $\lambda_{7}=0$ the non-zero rotation parameter is $t=ib$. The latter has significant impact on the spin coefficients. When $t=a$ the spin coefficients $\epsilon,\gamma,\alpha,\beta$ depends on the directional derivatives of real quantities yielding the following relations
@@
The second kind of simplification takes place when we substitute the four tilded spin relations into the CR and we compare the outcome with the NPE (f),(l),(e),(o) resulting in the key relations. The key relations help us to unfold the branches of the solutions. We postulate the most general case of the obtained relations after the comparison with NPE.
@@
Remark 1.
The application of null rotation and Lorentz transformations are transformative processes that yield valuable relationships connecting not only the spin coefficients amongst themselves but also with the Weyl components through the commutation relations, once the tilded spin coefficients have been annihilated. The outcome critically hinges on the preservation of the structure of the Killing tensor and the form of the spin coefficients. Importantly, the absence of either $\lambda_{0}$ or $\lambda_{7}$ is the catalyst for these relationships. Employing the complete form of the Killing tensor $K^{1,2,3}_{\mu\nu}$, conversely, provides no insight into this implied symmetry since both $t$ and $p$ are nullified.
@@ 4 The canonical forms of Killing tensor
The preservation of geometry of spacetime through a transformation reveals the existence of symmetries. These symmetries leave invariant the elements which characterize the geometry of the manifold, the metric tensor and the action apparently. Thus, the consideration of symmetries in the resolution process of the Einstein’s equations is indispensable since the non-linear character of the equations of gravity obligates us to do so, in order to obtain exact solutions of EFE.
In the following sections we present the definition of Killing tensor, the Killing equation their correlation to hidden symmetries. Afterwards, we demonstrate the order of arguments that were used by Churchill resulting in to the classification of the canonical forms of the components of a symmetric linear vector function or equivalently with an abstract symmetric tensor of valence 2.
@@ 4.1 Killing Tensor
where $p_{\mu}$ are the coordinates on the cotangent spaces or equivalently the canonical momenta of an observer. Then, an integral of motion could be defined as follows [25], [26].
@@
Killing tensors of rank r give rise to a homogeneous constant of motion of degree r in momenta. The inhomogeneous polynomial integrals of geodesic motion can be decomposed to their homogeneous parts and also to the corresponding parts that are associated with the Killing tensors with the equivalent rank. The Killing tensor of rank 1 equals to Killing vector and generates continuous symmetry transformations, these symmetries are called explicit [28]. The symmetries that correspond to higher-order ranks of momenta associated with Killing tensors of rank ($r>1$) are called hidden symmetries [29]. The investigation of special objects such Killing tensor or Killing-Yano is basically a devilish way to peep into the phase space searching for hidden symmetries [30]. Indeed, the explicit symmetries in a Hamiltonian system always could be “dragged up” instead of hidden symmetries. These kinds of symmetries emerge during the study of the dynamics of a system featuring the conserved quantities of the system or one-parameter isometries which is equivalent with the admission of existence of Killing vectors.
As Eisenhart [11] and Kalnins-Miller [12], [13], [14] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of order two [15]. Indeed, there is a bizarre relation between the structure of separated metrics with the structure of its characteristic Killing tensor. Benenti and Francaviglia [31] present a certain example where the additional information about the metric tensor serve as a catalyst in order to obtain the structure of the Killing tensor.
After this brief review about the Killing tensor let us proceed with its definition: Any symmetric tensor of order 2 whose the symmetric part of his covariant derivative vanishes is called a Killing tensor.